Part A Part complete ΔH∘rxn=+ 83 kJ , ΔSrxn=+ 146 J/K , T= 309 K Express...

Part A

Part complete

ΔH∘rxn=+ 83 kJ , ΔSrxn=+ 146 J/K , T= 309 K

Express your answer as an integer.

--

Part B

Part complete

ΔH∘rxn=+ 83 kJ , ΔSrxn=+ 146 J/K , T= 759 K

Express your answer as an integer.

--

ΔG

Δ G

=
kJ

Part C

Part complete

ΔH∘rxn=+ 83 kJ , ΔSrxn=− 146 J/K , T= 309 K

Express your answer as an integer.

--

ΔG

Δ G

=

nothing

kJ

Part D

Part complete

ΔH∘rxn=− 83 kJ , ΔSrxn=+ 146 J/K , T= 407 K

Express your answer as an integer.

--

ΔG

Δ G

=

nothing

kJ

SubmitPrevious AnswersRequest Answer

Part E

Complete previous part(s)

Part F

Complete previous part(s)

Part G

Complete previous part(s)
ΔG

Δ G

=

nothing

kJ
Answer #1

Gibb's free energy (G) is a thermodynamic quantity which is an indication of maximum amount of useful work that can be done.

Mathematically it is expressed as

G = H - TS

Hence change in Gibb's free energy is

G = H - (TS)

If Temperature is constant then

G = H - TS

PART A:

H = 83 KJ = 83000 J

S = 146 J/K , T = 309 K

G = 83000 - 146*309 = 37886 J = 37.886 KJ = 38 KJ (ROUND FIGURE TO INTEGER)

PART B:

H = 83 KJ = 83000 J

S = 146 J/K , T = 759 K

G = 83000 - 146*759 = -27814 J = -27.814 KJ = - 28 KJ

PART C:

H = 83 KJ = 83000 J

S = -146 J/K , T = 309 K

G = 83000 - (-146)*309 = 128114 J = 128.114 KJ = 128 KJ

PART D:

H = - 83 KJ = - 83000 J

S = 146 J/K , T = 407 K

G = - 83000 - 146*407 = -142422 J = -142.422 KJ = -142 KJ

Please do leave me your feedback. I'll be grateful for that :)